# Hallucinate at the Last in Long Response Generation: A Case Study on Long Document Summarization

### v1

@@ 4.1 Is it Intrinsic to Summarization?
This hypothesis is based on the intrinsic structure of human-written and model-generated summaries. In most summarization tasks—especially in news, scientific, and Wikipedia-style documents—the main ideas and most salient information are typically presented early in the summary Hermann et al. (2015); Narayan et al. (2018); Cohan et al. (2018). This structure mirrors the lead bias often observed in source documents, where the beginning contains the most crucial content Kim et al. (2019); Zhao et al. (2022); Ravaut et al. (2024). Motivated by this, our study investigates whether such structural bias also manifests in the form of decreasing factual consistency in generated outputs. Specifically, we compare the positional faithfulness scores of human-written reference summaries and model-generated summaries to assess whether the decline in faithfulness is an artifact of summarization’s inherent structure or a model-specific behavior.
Figure 5: Comparison of faithfulness scores between the reference summary and summaries generated by the Llama3.1-8B-Instruct model across increasing output lengths on the CNNDM dataset.
@@
We use human-written reference summaries from the CNNDM dataset, selecting examples with long summaries and input contexts of around 2K tokens. We compare these references to summaries generated by the Llama3.1-8B-Instruct model, evaluated using the metric described in Section 3.1.
@@
Figure 5 presents a positional faithfulness comparison between human-written reference summaries and model-generated summaries from the CNNDM dataset, analyzed across increasing output lengths. As shown in the results, the reference summaries exhibit a different pattern compared to the generated summaries. While the faithfulness scores of the generated summaries drop below 0.6 in the final segment, the reference summaries show a temporary dip around the middle but recover toward the end. This observation suggests that the Hallucination at the Last phenomenon cannot be solely attributed to the inherent characteristics of summarization.
(a) Average attention weights on the first, middle, and last sentences of summaries generated by the Llama3.1-8B-Instruct model.
(b) Average attention weights on the first, middle, and last sentences of summaries generated by the Mistral-7B-Instruct-v0.3 model.
@@ 4.2 Is it Intrinsic to Attention?
Unlike previous studies that focus on attention to the first generated token Hsieh et al. (2024), we compute attention weights at the sentence level. Moreover, we segment the full sequence into chunks of 100 tokens. Our analysis specifically focuses on three positions within the generated output: the first sentence, the middle sentence, and the final sentence. See Appendix E for more details.
@@
Figure 6 presents a visualization of average attention weights across sentences. As shown in the results, the Llama3.1-8B-Instruct model—which exhibited the Hallucinate at the Last pattern in Figure 3—assigns nearly four times more attention to the final sentence of the generated summary compared to the first and middle sentences. In contrast, the Mistral-7B-Instruct-v0.3 model assigns similar levels of attention to all three sentence positions. These findings suggest that increased attention to previously generated text correlates with a higher likelihood of hallucination. Moreover, we observe that for most models—excluding Mistral—as output length increases, attention becomes increasingly concentrated on generated tokens, which in turn amplifies hallucination.
5 How to Mitigate the Hallucination at the Last?
We investigate how we can resolve the Hallucinate at the Last phenomenon. To this end, we apply four methods.
@@ 5 How to Mitigate the Hallucination at the Last?
In this experiment, we generate summaries for the Wikipedia dataset using four different methods, with a context length of 7K tokens. As a baseline, we use summaries generated by the Llama3.1-8B-Instruct model. For comparison, we evaluate the following four methods:
@@
  • AdaCAD Wang et al. (2025) enhances factual consistency during generation by context-aware decoding.
See Appendix F for more experimental details.
@@
Table 4: Comparison of faithfulness scores and Sensitivity for long summaries generated by different mitigation methods across output bins on the Wikipedia dataset with a 7K context length.
@@
We report the faithfulness scores across output bins and the corresponding Sensitivity results in Table 4, and the faithfulness scores across increasing output lengths for each method in Figure 7. As shown in Table 4, BooookScore achieves the lowest sensitivity (2.0), making it the closest to zero among the four methods. Moreover, as illustrated in Figure 7, BooookScore maintains the highest level of faithfulness, particularly around the 30th sentence, and notably avoids the sharp decline in faithfulness observed in other methods at the end of the summary. These results suggest that generating summaries independently for each chunk and subsequently merging them can be an effective strategy for mitigating the Hallucinate at the Last phenomenon. However, while BooookScore mitigates the phenomenon, it doesn’t fundamentally solve the problem of generating long responses directly from raw long inputs, relying instead on decomposing the problem via chunking and merging. This highlights the necessity of future work to develop methods that enable models to maintain faithfulness when generating long text directly from long contexts.
@@ 6 Related Work
Recently, several studies have begun to explore long-output generation in LLMs Wu et al. (2025b); Ye et al. (2025). However, these works focus on procedural generation tasks that are not grounded in contextual input, limiting their applicability to real-world summarization scenarios. A notable study on multi-document summarization aligns with our findings Belém et al. (2025). Nevertheless, it does not address positional context-aware faithfulness over extended summary lengths.
@@
Figure 7: Comparison of faithfulness scores for long summaries generated by different mitigation methods across increasing output lengths, using the Wikipedia dataset with a 7K context length.
@@ 7 Conclusion
We identified and characterized the Hallucinate at the Last phenomenon in LLM-based long response generation, specifically in long document summarization. Our findings show that hallucinations disproportionately increase towards the end of long outputs, a bias amplified in longer summaries. We investigated the contributing factors and explored targeted mitigation strategies. Our work highlighted the importance of the output’s temporal dimension in LLM faithfulness and motivates future research into spatially-aware generation techniques.
@@
In this study, we explored four domains, primarily because the corresponding datasets provided input contexts of the desired length. However, we were not able to investigate domains such as books, dialogues, movie scripts, or meeting transcripts. Books were excluded due to their excessive length, while the other domains lacked datasets with suitable input lengths or sufficient sample sizes. Additionally, we also did not explore models of varying sizes.
@@
Ethics Statement
This study leverages publicly available datasets, including Wikipedia, Arxiv, Pubmed, Govreport, and CNNDM, to analyze long-form text generation in LLMs. All experiments were conducted in a consistent and reproducible manner across models and datasets, without any manipulation or omission of data or results. We investigate the phenomenon of Hallucinate at the Last, a tendency observed in long context-aware summarization models to generate hallucinated content toward the final portion of the summary. By drawing attention to this issue, our study contributes to ongoing efforts to enhance the reliability and factual consistency of LLM-generated summaries.
@@ Acknowledgement
This work was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) [RS-2021-II211341, Artificial Intelligent Graduate School Program (Chung-Ang University)].
@@ Appendix A Details for Generating Summaries
Short outputs were standardized to a length of 100 to 200 words, while long outputs were approximated by converting 30% of the total context token length into a corresponding word range. The specific word ranges used for prompting long-summary generation are provided in Table 5, and the prompt template employed for summary generation is shown in Table 7.
@@
Table 6 presents the average word counts of both short and long summaries generated by each model. As the results indicate, LongWriter-Llama3.1-8B frequently produces considerably longer outputs, while Qwen2.5-7B-Instruct tends to generate relatively shorter summaries.
@@
short long
Llama3.1-8B-Instruct 279 834
Qwen2.5-7B-Instruct 153 751
Mistral-7B-Instruct-v0.3 285 794
GPT-4o mini 172 925
LongWriter-Llama3.1-8B 398 3291
Table 6: The average word counts of both short and long summaries generated by each model on the Wikipedia 7K dataset.
Appendix B Details for Evaluation Metric
In this section, we present the experimental details of the evaluation metric used to compute sentence-level faithfulness scores in Section 3.1.
For atomic fact decomposition, we utilize the Llama3.1-8B-Instruct model with vLLM Kwon et al. (2023) [^1] https://github.com/vllm-project/vllm. For the NLI model, we use the state-of-the-art hallucination evaluation model Bao et al. (2024) [^2] https://huggingface.co/vectara/hallucination_evaluation_model.
Appendix C More Results on Varying Models
Usage of LLMs
We utilized publicly available instruction-tuned models, including Llama3.1-8B-Instruct [^3] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct, Mistral-7B-Instruct-v0.3444https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3, and Qwen2.5-7B-Instruct [^5] https://huggingface.co/Qwen/Qwen2.5-7B-Instruct from HuggingFace. For the GPT-4o mini model, we used gpt-4o-mini-2024-07-18. All summaries were generated using greedy decoding with float16 precision.
More Results
We report the faithfulness scores and Sensitivity for short summaries generated by different models across output bins on the Wikipedia dataset in Table 8, and those of long summaries in Table 9.
@@ Appendix D More Results on Varying Datasets
We report comparison of faithfulness scores in short summaries generated by the Llama 3.1 model at various output positions across multiple domains, with context lengths ranging from 1K to 8K in Table 8.
We report the faithfulness scores and Sensitivity of short summaries generated by the Llama3.1-8B-Instruct model across output bins in multiple domains in Table 10, and those of long summaries in Table 11.
As demonstrated by the experimental results, none of the summaries generated by the Llama3.1-8B-Instruct model across various domains exhibited a sensitivity value below zero.
@@ Appendix E Attention Weight Calculation
This section provides a more detailed and formal explanation our attention weight computation method introduced in Section 4.2.
@@
$x^{first}$ =$=\{x^{first}_{j}\}_{j=1}^{T_{1}}$ (6)
$x^{second}$ =$=\{x^{second}_{j}\}_{j=1}^{T_{2}}$
$x^{third}$ =$=\{x^{third}_{j}\}_{j=1}^{T_{3}}$

### v2

@@ 4.1 Is it Intrinsic to Summarization?
This hypothesis draws on the intrinsic structure of both human- and model-generated summaries. In typical summarization domains such as news, scientific articles, and Wikipedia-style documents, key information is predominantly introduced early in the summary Hermann et al. (2015); Narayan et al. (2018); Cohan et al. (2018), reflecting the lead bias observed in source texts, where the opening contains the most salient content Kim et al. (2019); Zhao et al. (2022); Ravaut et al. (2024). Building on this, our study investigates whether this structural bias is also reflected in decreasing factual consistency within generated outputs by comparing positional faithfulness scores between human-written and model-generated summaries.
Figure 5: Faithfulness scores of reference summaries and those generated by Llama on the CNNDM dataset.
@@
We use human-written reference summaries from the CNNDM dataset, selecting examples with long summaries and input contexts of around 2K tokens. We compare these references to summaries generated by the Llama3.1-8B-Instruct model.
@@
Figure 5 compares positional faithfulness between human-written and model-generated CNNDM summaries. We observe that the faithfulness of the generated summaries declines steadily, falling below 0.6 in the final segment, whereas reference summaries dip around the middle but recover. This suggests that Hallucination at the Last is not solely due to summarization’s inherent structure.
@@ 4.2 Is it Intrinsic to Attention?
Unlike previous studies that focus on attention to the first generated token Hsieh et al. (2024), we compute attention weights at the sentence level. Moreover, we segment the full sequence into chunks of 100 tokens. Our analysis specifically focuses on three positions within the generated output: the first sentence, the middle sentence, and the final sentence. Details of the attention weight calculation are provided in Appendix E.
@@
Figure 6 presents a visualization of average attention weights across sentences. As shown in the results, the Llama model, which exhibited the Hallucinate at the Last pattern in Figure 3, assigns nearly three times more attention to the final sentence of the generated summary compared to the first and middle sentences. In contrast, the Qwen model assigns similar levels of attention to all three sentence positions. These findings suggest that increased attention to previously generated text correlates with a higher likelihood of hallucination.
This consistent distribution in Qwen can be attributed to its sliding window attention mechanism Child et al. (2019), which maintains a balanced focus across the context and mitigates the downstream attention concentration that triggers hallucinations. Table 5 compares the Qwen model with and without sliding window attention. In this setting, the sliding window context length is fixed at 16000 tokens. The results clearly show that sliding window attention omission increases sensitivity across all context and summary lengths, up to 9.75. They demonstrate that employing sliding window attention not only enhances decoding efficiency but also reduces attention concentration on later tokens as the generated sequence length increases.
(a) Average attention weights of summaries generated by Llama.
(b) Average attention weights of summaries generated by Qwen.
5 How to Mitigate the Hallucinate at the Last?
In the preceding section, we demonstrated that the Hallucinate at the Last phenomenon could be partially mitigated through the use of sliding window attention. In the following analysis, we investigate strategies to address this issue in models that are not trained with sliding window attention. To this end, we apply four methods that may mitigate the phenomenon.
Models Context Generated Summary Bins Sensitivity
Length 1 2 3 4 5
Standard Summary
Qwen2.5-7B-Instruct 1K 0.89 0.78 0.71 0.77 0.85 -6.25
2K 0.90 0.89 0.83 0.78 0.84 1.0
3K 0.87 0.84 0.79 0.83 0.82 1.25
4K 0.91 0.88 0.90 0.79 0.89 -2.0
5K 0.86 0.89 0.85 0.76 0.89 -5.0
6K 0.93 0.91 0.92 0.80 0.90 -1.0
7K 0.90 0.89 0.89 0.84 0.82 6.0
8K 0.89 0.93 0.88 0.89 0.92 -2.25
Without 1K 0.88 0.79 0.74 0.76 0.84 -4.75(+1.5)
2K 0.90 0.89 0.83 0.78 0.84 1.0(+0.0)
Sliding 3K 0.88 0.86 0.82 0.85 0.83 2.25(+1.0)
4K 0.91 0.88 0.87 0.83 0.83 4.25(+6.25)
Winodw 5K 0.87 0.87 0.81 0.82 0.80 4.25(+9.25)
6K 0.93 0.90 0.91 0.85 0.86 3.75(+4.75)
Attention 7K 0.91 0.91 0.89 0.84 0.79 9.75(+3.75)
8K 0.88 0.93 0.89 0.89 0.92 -2.25(+0.0)
Long Summary
Qwen2.5-7B-Instruct 1K 0.88 0.76 0.81 0.83 0.83 -1.0
2K 0.83 0.89 0.84 0.83 0.82 2.75
3K 0.83 0.84 0.86 0.79 0.84 -1.0
4K 0.87 0.83 0.79 0.83 0.84 -1.0
5K 0.87 0.89 0.89 0.80 0.85 1.25
6K 0.89 0.81 0.78 0.78 0.82 -0.5
7K 0.86 0.78 0.79 0.83 0.84 -2.5
8K 0.86 0.85 0.85 0.85 0.85 0.25
Without 1K 0.89 0.77 0.83 0.84 0.80 3.25(+4.25)
2K 0.83 0.89 0.83 0.80 0.78 5.75(+3.0)
Sliding 3K 0.83 0.84 0.86 0.79 0.84 -1.0(+0.0)
4K 0.87 0.83 0.79 0.83 0.84 -1.0(+0.0)
Winodw 5K 0.86 0.90 0.83 0.79 0.80 4.5(+3.25)
6K 0.89 0.82 0.84 0.83 0.78 6.5(+7.0)
Attention 7K 0.86 0.80 0.78 0.83 0.82 -0.25(+2.25)
8K 0.88 0.85 0.85 0.86 0.82 4.0(+3.75)
Table 5: Faithfulness scores and sensitivity for standard and long summaries generated by Qwen2.5-7B-Instruct on the Wikipedia dataset, with and without sliding window attention. Note that higher sensitivity indicates stronger hallucination effects toward the end of the output.
@@ 5 How to Mitigate the Hallucinate at the Last?
In this experiment, we generate Wikipedia summaries with a 7K context using four methods, taking Llama3.1-8B-Instruct as the baseline. For comparison, we evaluate the following four methods: (See Appendix F for more experimental details.)
@@
  • AdaCAD Wang et al. (2025) enhances factual consistency during generation by context-aware decoding Shi et al. (2024).
@@
Table 6: Faithfulness scores and sensitivity for long summaries generated by different mitigation methods on the Wikipedia dataset.
@@
We report the faithfulness scores across output bins and the sensitivity results in Table 6, and the faithfulness scores across increasing output lengths for each method in Figure 7. As shown in Table 6, BooookScore achieves the lowest sensitivity (2.0), making it the closest to zero among the four methods. Moreover, as illustrated in Figure 7, BooookScore maintains the highest level of faithfulness, particularly around the 30th sentence, and notably avoids the sharp decline in faithfulness observed in other methods at the end of the summary. These results suggest that generating summaries independently for each chunk and subsequently merging them can be an effective strategy for mitigating the Hallucinate at the Last phenomenon. Nevertheless, while BooookScore serves as a practical mitigation for models without sliding window attention, it does not fundamentally resolve the issue, as it relies on chunking and merging rather than direct long-context generation. This highlights the need for future work to develop methods that maintain faithfulness when generating long text directly from long contexts.
@@ 6 Related Work
Recently, several studies have begun to explore long-output generation in LLMs Wu et al. (2025b); Ye et al. (2025). However, these works focus on procedural generation tasks that are not grounded in contextual input, limiting their applicability to real-world summarization scenarios. A notable study on multi-document summarization aligns with our findings Belém et al. (2025); Zhao et al. (2025). Nevertheless, it does not address positional context-aware faithfulness over extended summary lengths.
@@
Figure 7: Faithfulness scores for long summaries generated by different mitigation methods on the Wikipedia dataset across increasing output lengths.
@@ 7 Conclusion
We identified and characterized the Hallucinate at the Last phenomenon in LLM-based long response generation, specifically in long document summarization. Our findings show that hallucinations disproportionately increase towards the end of long outputs, a bias amplified in longer summaries. We investigated the contributing factors and explored targeted mitigation strategies. Our work highlighted the importance of the output’s temporal dimension in LLM faithfulness and motivates future research into positionally-aware generation techniques.
@@
In this study, we explored four domains, primarily because the corresponding datasets provided input contexts of the desired length. However, we were not able to investigate domains such as books, dialogues, movie scripts, or meeting transcripts. Books were excluded due to their excessive length, while the other domains lacked datasets with suitable input lengths or sufficient sample sizes.
@@ Acknowledgement
This work was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) [RS-2021-II211341, Artificial Intelligence Graduate School Program (Chung-Ang University)].
@@ Appendix A Details for Generating Summaries
Standard outputs were standardized to a length of 100 to 200 words, while long outputs were approximated by converting 30% of the total context token length into a corresponding word range. The specific word ranges used for prompting long-summary generation are provided in Table 7, and the prompt template employed for summary generation is shown in Table 10.
Context Length Words Range
1K 200 to 250
2K 400 to 500
3K 600 to 750
4K 800 to 1000
5K 1000 to 1250
6K 1200 to 1500
7K 1400 to 1750
8K 1600 to 2000
Table 7: Words range used in the prompt for long output generation.
@@
Table 8 presents the average word counts of both standard and long summaries generated by each model. As the results indicate, LongWriter-Llama3.1-8B frequently produces considerably longer outputs, while Qwen2.5-7B-Instruct tends to generate relatively shorter summaries.
@@
standard long
[…]
Table 8: The average word counts of both standard and long summaries generated by each model on the Wikipedia 7K dataset.
Appendix B Details for Model Usage
In this section, we present the experimental details for the model usage in Section 3.
Details for Evaluation Method
For atomic fact decomposition, we utilize the Llama3.1-8B-Instruct model with vLLM Kwon et al. (2023) [^2] https://github.com/vllm-project/vllm. To further illustrate the evaluation process, we present in Table 9 a concrete example that demonstrates how a given sentence is decomposed into its constituent atomic facts and subsequently assessed for faithfulness. This example clarifies the step-by-step procedure and provides transparency regarding the alignment between the generated atomic facts and the underlying reference information.
First Sentence
Sara Sorribes Tormo is a talented Spanish professional tennis player born on October 8, 1996, in Castellón de la Plana, Spain.
Atomic Facts Score
Sara Sorribes Tormo is a Spanish professional tennis player. 0.92
Sara Sorribes Tormo was born in Spain. 0.90
Sara Sorribes Tormo was born on October 8. 0.87
Sara Sorribes Tormo was born in 1996. 0.96
Sara Sorribes Tormo was born in Castellón de la Plana. 0.93
Avg. Score 0.92
Last Sentence
As Sara navigates her career, her unique blend of skill, perseverance, and spirit positions her as a promising athlete, inspiring both current and aspiring players.
Atomic Facts Score
Sara navigates her career. 0.87
Sara’s career is that of an athlete. 0.91
Sara has a unique blend of skill. 0.14
Sara has perseverance. 0.81
Sara is a promising athlete. 0.30
Sara inspires current players. 0.11
Sara inspires aspiring players. 0.15
Avg. Score 0.47
Table 9: Example of the evaluation procedure with the MiniCheck-Flan-T5-Large Model.
Details for LLMs Usage
We utilized publicly available instruction-tuned models, including Llama3.1-8B-Instruct [^3] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct, Gemma3-12B-Instruct [^4] https://huggingface.co/google/gemma-3-12b-it, and Qwen2.5-7B-Instruct [^5] https://huggingface.co/Qwen/Qwen2.5-7B-Instruct from HuggingFace. For the GPT-4o mini model, we used gpt-4o-mini-2024-07-18. All summaries were generated using greedy decoding with float16 precision.
Appendix C More Results with Different Methods
We adopt a diverse set of methods to validate the robustness of our faithfulness evaluation framework. In this section, we present experimental results for faithfulness assessment models beyond MiniCheck, incorporating approaches based on NLI, ROUGE-L, and LLMs. Specifically, we substitute the scoring model $\mathcal{M}$ in Equation 2 of Section 3 with these alternative methods.
Minicheck Based
We utilize MiniCheck-Flan-T5-Large Model [^6] https://huggingface.co/lytang/MiniCheck-Flan-T5-Large for the main evaluation method. We report the faithfulness scores and sensitivity for standard summaries generated by different models across output bins on the Wikipedia dataset in Table 11, and those of long summaries in Table 12.
NLI Based
For the NLI model, we use the state-of-the-art hallucination evaluation model Bao et al. (2024) [^7] https://huggingface.co/vectara/hallucination_evaluation_model.
We report the faithfulness scores and sensitivity assessed by NLI model for standard summaries generated by different models across output bins on the Wikipedia dataset in Table 13, and those of long summaries in Table 14. As observed in the experimental results, consistent with the main results, the Llama and GPT models consistently exhibit high sensitivity, whereas the Qwen model occasionally demonstrates sensitivity values even below zero.
ROUGE Based
We report the faithfulness scores and sensitivity assessed by ROUGE-L Lin (2004) for standard summaries generated by different models across output bins on the Wikipedia dataset in Table 17, and those of long summaries in Table 18.
LLM Based
We present the error rates assessed by FineSurE framework for standard summaries generated by different models across output bins on the Wikipedia dataset in Table 15, and those of long summaries in Table 16.
As observed in the experimental results, consistent with the main results, the Llama and GPT models consistently exhibit the highest error rates in the final bin. In contrast, the Qwen model does not yield the highest error rates in the final bin for both the cases of standard and long summaries.
@@ Appendix D More Results on Varying Datasets
We report comparison of faithfulness scores in standard summaries generated by the Llama model at various output positions across multiple domains, with context lengths ranging from 1K to 8K in Figure 8.
Minicheck Based
We report the faithfulness scores and sensitivity assessed by MiniCheck-Flan-T5-Large model of standard summaries generated by the Llama3.1-8B-Instruct model across output bins in multiple domains in Table 19, and those of long summaries in Table 20.
The experimental results reveal that in the case of standard summaries, sensitivities below zero are frequently observed; notably, in the Pubmed dataset, more than half of the values fall below zero. In contrast, such negative sensitivities are rarely observed in long summaries, with the Pubmed dataset in particular showing no sensitivities below zero, unlike the standard summary setting.
NLI Based
We report the faithfulness scores and sensitivity assessed by NLI model for standard summaries generated by the Llama3.1-8B-Instruct model across output bins in multiple domains in Table 21, and those of long summaries in Table 22.
As demonstrated by the experimental results, consistent with the main results, none of the summaries generated by the Llama3.1-8B-Instruct model across various domains exhibited a sensitivity value below zero.
@@ Appendix E Attention Weight Calculation
This section provides a more detailed and formal explanation of our attention weight computation method introduced in Section 4.
@@
$x^{first}$ $=\{x^{first}_{j}\}_{j=1}^{T_{1}}$ (6)
$x^{second}$ $=\{x^{second}_{j}\}_{j=1}^{T_{2}}$
$x^{third}$ $=\{x^{third}_{j}\}_{j=1}^{T_{3}}$